OpenStudy (anonymous):

The highest elevation on earth is Mount Everest in Asia. It is approximately 2.9 x 104 feet high. If a climber only made it half way up the mountain, how high did he climb? Write the answer in scientific notation.

OpenStudy (campbell_st):

its simply \[(2.9 \times10^4) \div 2 =\]

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

tell me the answer

OpenStudy (campbell_st):

here is a scientific calculator you can use to get the answer http://web2.0calc.com/

OpenStudy (campbell_st):

its the same as \[29000\div 2 =\]

OpenStudy (anonymous):

\[[2.9*10^{4}]/2=\frac{29000}{2}=14500\] do you know how to write 14500 in scientific notation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ah, there lies the problem. Want me to teach you?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

lol... why do the question...

OpenStudy (anonymous):

um...

OpenStudy (anonymous):

If you have something like 156900, you just put a point after the first non-zero digit like this: 1.56900 Now what should you multiply 1.56900 with to get 156900? 100000, right? That is 10^5. \[\therefore 156900=1.569*10^{5}\]

OpenStudy (anonymous):

thanx your the best